OpenStudy (anonymous):

Am I right or wrong? f(x)=-(x+1)^2+4 vertex: (-1,4) x= .4 x= -2.4 y= 5

OpenStudy (anonymous):

@Hero

hero (hero):

I agree with your vertex. What does the rest of the information represent. I assume the rest of what you posted are the x and y intercepts.

OpenStudy (anonymous):

Yes they're. But, I'm not entirely positive I got them properly, I can type up the equation, if you'd like.

OpenStudy (jdoe0001):

\(\bf y = -(x+1)^2+4\implies 0=-x^2-2x-1+4\implies 0=-x^2-2x+3\\ \quad \\ \textit{using the quadratic formula}\\ \quad \\ x= \cfrac{ - (-2) \pm \sqrt { (-2)^2 -4(-1)(3)}}{2(-1)}\implies x= \cfrac{ 2 \pm \sqrt { 4+12}}{-2}\)

hero (hero):

Here's the graph of it: https://www.desmos.com/calculator/lg1z24reec

OpenStudy (anonymous):

pellet, it was that formula?! I was using the (a+h)^2+k

hero (hero):

There's no reason to use the quadratic formula to find the x and y intercepts.

OpenStudy (anonymous):

I'll fix it up now. Dx

OpenStudy (anonymous):

Wait, I don't understand... Which formula then?

hero (hero):

Instead, what you do is: First remember y = f(x) To find the x intercept, set y = 0, then solve for x: y=-(x+1)^2+4 0 = -(x+1)^2 + 4 (x+1)^2 = 4 x + 1 = ±2 x = ±2 - 1 `x = 1` `x = -3` To find the y-intercept, set x = 0, then solve for y y = -(0+ 1)^2 + 4 y = -(1)^2 + 4 y = -1 + 4 `y = 3`

OpenStudy (anonymous):

I see where I messed up now. Thank you!

OpenStudy (anonymous):

Okay, let me make sure that I'm absolutely doing it right. f(x)=(x+4)^2-2 vertex= (-4,-2) 0=(x+4)^2-2 2=(x+4)^2 1.4= x+4 -1.4-4 = -5.4 1.4-4=-2.6 f(0)=(0+4)^2-2 4^2-2 16-2 14 Is that done right?'' @Hero

hero (hero):

Correct

OpenStudy (anonymous):

Yay!! Okay, now these two are different but I'm positive they're right. f(x)=-x^2+2x+3 -(2/2(-1)) = -2/-2 = 1 0=-x^2+2x+3 0=(x-1)(x+3) x=1 x=-3 y=c so... y=3 f(1)=1^2+2+3 1+2+3 y=5 (1,5) Sorry for being bothersome, by the way;;

OpenStudy (anonymous):

and then there's f(x)=2x^2-4x-6 -(-4/2(2)= 4/4 = 1 f(1)= 2^2-4-6 4-4-6 0-6 -6 (1,-6) f(0)=2(0)^2-4(0)-6 0-0-6 y=-6 0=2x^2-4x-6 2(x^2-2x-3) 2(x+1)(x-3) x= 1.4 x= -1 x= 3

OpenStudy (anonymous):

@Hero

hero (hero):

I'm about to go take a nap. ttyl

OpenStudy (anonymous):

@phi , do you think you can help me? :c

OpenStudy (phi):

*** 0=-x^2+2x+3 0=(x-1)(x+3) x=1 x=-3 *** the minus out front changes things. I would factor it out to get -( x^2 -2x -3) = 0 (x-3)(x+1)=0 x=3 , x= -1

OpenStudy (anonymous):

Yeah--- I see where I got it wrong... But other than that, the rest is fine, right? c:

OpenStudy (phi):

*** f(x)=2x^2-4x-6 -(-4/2(2)= 4/4 = 1 f(1)= 2^2-4-6 *** x value of the vertex is 1. the y value is 2*1^2 - 4*1 -6 = 2 - 4 - 6= 2-10= -8

OpenStudy (phi):

it looks like you have the right idea, but you have to be more careful when doing the work.

OpenStudy (anonymous):

Ahh-- Dang, I didn't realize that. But thank you so much!! I really appreciate it.

OpenStudy (anonymous):

Would you mind helping me with four other equations?

hero (hero):

@TheHero, when doing these, you can easily check your own work.